# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ Abstract
1 Introduction
Lyapunov theory is one of the most widely-used approaches to stability analysis of a nonlinear system Isidori [1995], Marino and Tomei [1995], Khalil [2002], Vidyasagar [2002], Slotine and Li [1991], Nijmeijer and Van der Schaft [1990], which provides a condition for stability with respect to an equilibrium point, a target trajectory, or an invariant set. Contraction theory Lohmiller and Slotine [1998, 2000], Slotine and Lohmiller [2001], Slotine [2003], Wang and Slotine [2005] rewrites suitable Lyapunov stability conditions using a quadratic Lyapunov function of the differential states, defined by a Riemannian contraction metric and its uniformly positive definite matrix, thereby characterizing a necessary and sufficient condition for incremental exponential convergence of the multiple nonlinear system trajectories to one single trajectory. It can be regarded as a generalization of Krasovskii’s theorem [Slotine and Li, 1991, p. 83] applied to nonlinear incremental stability analysis Lohmiller and Slotine [1998], Angeli [2002], where the differential formulation permits a pure differential coordinate change with a non-constant metric for simplifying its stability proofs Lohmiller and Slotine [1998].
The differential nature of contraction theory implies we can exploit the Linear Time-Varying (LTV) systems-type techniques for nonlinear stability analysis and control/estimation synthesis Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Dani et al. [2015], Tsukamoto and Chung [2019] (see Table 1). We emphasize that some of these methodological simplifications in contraction theory are accomplished by its extensive use of exponential stability along with the comparison lemma [Khalil, 2002, pp. 102-103, pp. 350-353], in lieu of Input-to-State Stability (ISS) or uniform asymptotic stability which often renders nonlinear stability analysis more involved Isidori [1995], Marino and Tomei [1995], Khalil [2002], Vidyasagar [2002], Slotine and Li [1991], Nijmeijer and Van der Schaft [1990]. Several studies related to the notion of contraction, although not based on direct differential analysis, can be traced back to Hartman [2002], Demidovich [1961], Lewis [1949].
@@ 1 Introduction
7. $\mathcal{L}_{2}$ stability condition Reduces to LMI conditions in terms of contraction metric defined by positive definite matrix $M$ Hamilton-Jacobi inequality (PDE) in terms of Lyapunov function $V$ [Khalil, 2002, p. 211]
8. Modular stability Differential analysis handles hierarchical, feedback, and parallel combinations Slotine [2003] Passivity is not intuitive for hierarchical combinations
@@ 1.1 Paper Organization
In Sec. 2, we present fundamental results of contraction theory for nonlinear robustness and stability analysis. Sections 3 and 4 consider nonlinear optimal feedback control and estimation problems from the perspective of contraction theory, deriving and delineating a convex optimization-based method for constructing contraction metrics. Section 3 also presents some new results on relating contraction theory to the bounded real lemma Boyd et al. [1994] and Kalman–Yakubovich–Popov (KYP) lemma [Dullerud and Paganini, 2000, p. 218].
@@ 1.2 Related Work
According to contraction theory, all the solution trajectories of a given nonlinear system converge to one single trajectory incrementally and exponentially, regardless of the initial conditions, if the system has a contraction metric and its associated quadratic Lyapunov function of the differential state Lohmiller and Slotine [1998] (see 1133 of Table 1). This paper primarily considers this generalized notion of stability, called incremental exponential stability, which enables systematic learning-based and data-driven control synthesis with formal robustness and stability guarantees. The purpose of this section is not for proposing that other notions of stability, such as traditional Lyapunov-based stability or incremental asymptotic stability for semi-contraction systems, should be replaced by incremental exponential stability, but for clarifying its advantages to help determine which of these concepts is the best fit when analyzing nonlinear robustness and stability (see Sec. 2 for the illustrative examples).
@@
In keeping with the use of the comparison lemma [Khalil, 2002, pp. 102-103, pp. 350-353], incremental exponential stability naturally holds for non-autonomous nonlinear systems without any additional conditions or modifications unlike Lyapunov techniques (see, e.g., the examples and theorems in Rantzer [2001]). Such aspects of contraction theory, including the extensive use of exponential stability, result in intuitive proofs on ISS and finite-gain $\mathcal{L}_{p}$ stability both for autonomous and non-autonomous nonlinear systems, without resorting to uniform asymptotic stability which makes stability analysis much more involved than necessary Isidori [1995], Marino and Tomei [1995], Khalil [2002], Vidyasagar [2002], Slotine and Li [1991], Nijmeijer and Van der Schaft [1990]. In particular, perturbed systems with a time-varying target trajectory are non-autonomous, and thus contraction theory allows us to easily obtain an explicit exponential bound on its tracking error Lohmiller and Slotine [1998], Chung et al. [2013], Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019], leveraging incremental stability of the perturbed system trajectories with respect to the target trajectory (see $4$ of Table 1). Having such analytical bounds on the tracking error is almost essential for the safe and robust implementation of automatic control schemes in real-world scenarios.
@@
Contraction theory also simplifies input-output stability analysis, such as $\mathcal{L}_{p}$ gain analysis of nonlinear systems including the $\mathcal{H}_{\infty}$ nonlinear optimal control problem Van der Schaft [1992], Isidori and Astolfi [1992], Doyle et al. [1989], Khargonekar and Rotea [1991], Xie and de Souza Carlos [1992], Scherer et al. [1997], Hinrichsen and Pritchard [1998], Chen and Zhang [2004]. For example, in Lyapunov theory, the problem of finding a suitable Lyapunov function with the smallest $\mathcal{L}_{2}$ gain boils down to solving a Partial Differential Equation (PDE) called the Hamilton-Jacobi inequality [Khalil, 2002, p. 211]Van der Schaft [1992] in terms of its associated Lyapunov function. In essence, since contraction theory utilizes a quadratic Lyapunov function of the differential state for stability analysis, the problem could be solved with a Linear Matrix Inequality (LMI) constraint Boyd et al. [1994] analogous to the KYP lemma [Dullerud and Paganini, 2000, p. 218] in LTV systems theory Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019], Manchester and Slotine [2018], as shall be shown in Sec. 34 (see 5577 of Table 1). There exist stochastic analogues of these stability results for nonlinear systems with stochastic perturbations Pham et al. [2009], Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019], Han and Chung [2021], as shall be outlined also in this paper.
@@
Another notable feature of contraction theory is modularity, which preserves contraction through parallel, feedback, and hierarchical combinations Slotine [2003], Slotine and Lohmiller [2001], specific time-delayed feedback communications Wang and Slotine [2006], synchronized coupled oscillations Wang and Slotine [2005], Slotine et al. [2004], and synchronized networks Slotine [2003], Chung et al. [2013], Pham and Slotine [2007], Slotine et al. [2004], Russo et al. [2013], Chung et al. [2013], Chung and Slotine [2009], expanding the results obtainable with the passivity formalism of Lyapunov theory Popov and Georgescu [1973][Khalil, 2002, p. 227][Slotine and Li, 1991, p. 132] (see $8$ of Table 1). Due to all these useful properties, extensions of contraction theory have been considered in many different settings. These include, but are not limited to, stochastic contraction (Gaussian white noise Pham et al. [2009], Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019], Poisson shot noise and Lévy noise Han and Chung [2021]), contraction for discrete and hybrid nonlinear systems Lohmiller and Slotine [1998], Slotine et al. [2004], Lohmiller and Slotine [2000], Pham [2008], Tsukamoto and Chung [2021, 2019], Wang et al. [2021], partial contraction Wang and Slotine [2005], transverse contraction Manchester and Slotine [2014], incremental stability analysis of nonlinear estimation (the Extended Kalman Filter (EKF) Bonnabel and Slotine [2015], nonlinear observers Zhao and Slotine [2005], Dani et al. [2015], Simultaneous Localization And Mapping (SLAM) Tan et al. [2017]), generalized gradient descent based on geodesical convexity Wensing and Slotine [2020], contraction on Finsler and Riemannian manifolds Forni and Sepulchre [2014], Pham and Slotine [2013], Simpson-Porco and Bullo [2014], Chaffey and Manchester [2018], contraction on Banach and Hilbert spaces for PDEs Shafi et al. [2013], Bonnabel et al. [2011], Cisneros-Velarde et al. [2020], non-Euclidean contraction Aminzare and Sontag [2014], Davydov et al. [2021], contracting learning with piecewise-linear basis functions Lohmiller et al. [2018], incremental quadratic stability analysis Luis D’Alto [2013], contraction after small transients Margaliot et al. [2016], immersion and invariance stabilizing controller design Wang et al. [2017], Astolfi and Ortega [2003], and Lipschitz-bounded neural networks for robustness and stability guarantees Shi et al. [2019], Revay et al. [2020, 2021], Manchester et al. [2021].
@@
The benefits of contraction theory reviewed so far naturally lead to a discussion on how to design a contraction metric and corresponding Lyapunov function. There are some cases in which we can analytically find them using special structures of systems in question Isidori [1995], Slotine and Li [1991], Khalil [2002]. Among these are Lagrangian systems [Slotine and Li, 1991, p. 392], where one easy choice of positive definite matrices that define a contraction metric is the inertia matrix, or feedback linearizable systems Freeman and Kokotovic [1995], Primbs et al. [1999], Prajna et al. [2004], Aylward et al. [2008], Chesi [2010], where we could solve the Riccati equation for a contraction metric as in LTV systems. This is also the case in the context of state estimation (e.g., the nonlinear SLAM problem can be reformulated as an LTV estimation problem using virtual synthetic measurements Tan et al. [2017], Dani et al. [2015]). Once we find a contraction metric and Lyapunov function of a nominal nonlinear system for the sake of stability, they could be used as a Control Lyapunov Function (CLF) to attain stabilizing feedback control Sontag [1983, 1989], Khalil [2002] or could be augmented with an integral control law called adaptive backstepping to recursively design a Lyapunov function for strict- and output-feedback systems Krstic et al. [1995], Deng and Krstic [1997, 1999], Sharma and Kar [2009]. However, deriving an analytical form of contraction metrics for general nonlinear systems is challenging, and thus several search algorithms have been developed for finding them at least numerically using the LMI nature of the contraction condition.
@@
The simplest of these techniques is the method of State-Dependent Riccati Equation (SDRE) Cloutier [1997], Banks et al. [2007], Çimen [2012], which uses the State-Dependent Coefficient (SDC) parameterization (also known as extended linearization) of nonlinear systems for feedback control and state estimation synthesis. Motivating optimization-based approaches to design a contraction metric, it is proposed in Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019] that the Hamilton-Jacobi inequality for the finite-gain $\mathcal{L}_{2}$ stability condition can be expressed as an LMI when contraction theory is equipped with the extended linearity of the SDC formulation. Specifically, in Tsukamoto and Chung [2021, 2021], Tsukamoto et al. [2021], Tsukamoto and Chung [2019], a convex optimization-based framework for robust feedback control and state estimation, named ConVex optimization-based Steady-state Tracking Error Minimization (CV-STEM), is derived to find a contraction metric that minimizes an upper bound of the steady-state distance between perturbed and unperturbed system trajectories. In this context, we could utilize Control Contraction Metrics (CCMs) Manchester and Slotine [2017], Singh et al. [2017, 2020], Wang et al. [2019, 2020], Manchester and Slotine [2018] for extending contraction theory to the systematic design of differential feedback control $\delta u=k(x,\delta x,u,t)$ via convex optimization, achieving greater generality at the expense of computational efficiency in obtaining $u$. Applications of the CCM to estimation, adaptive control, and motion planning are discussed in Manchester [2018]Lopez and Slotine [2021, 2021], Boffi et al. [2020], and Singh et al. [2017], Lopez et al. [2019], Lakshmanan et al. [2020], Gahlawat et al. [2021], Sun et al. [2021], respectively, using geodesic distances between trajectories Pham and Slotine [2013]. It is also worth noting that the objective function of CV-STEM has the condition number of a positive definite matrix that defines a contraction metric as one of its arguments, rendering it applicable and effective even to machine learning-based automatic control frameworks as shall be seen in Sec. 59.
@@
One drawback of these numerical schemes is that they require solving optimization problems or nonlinear systems of equations at each time instant online, which is not necessarily realistic in practice. In Lyapunov theory, approximating functions in a given hypothesis space has therefore been a standard technique Johansson and Rantzer [1998], Johansen [2000], Giesl [2007], Papachristodoulou and Prajna [2002], Richards et al. [2018], Chang et al. [2019], Han and Panagou [2017], Wang and Jungers [2021], Chakrabarty et al. [2020a, b], Gaby et al. [2021], where examples of its function classes include piecewise quadratic functions Johansson and Rantzer [1998], linearly parameterized non-quadratic functions Johansen [2000], a linear combination of radial basis functions Giesl [2007], Sum-Of-Squares (SOS) functions Papachristodoulou and Prajna [2002], and neural networks Richards et al. [2018], Chang et al. [2019], Gaby et al. [2021]. In Aylward et al. [2008], Manchester and Slotine [2017], the SOS approximation is investigated for the case of contraction theory, showing that the contraction condition can be relaxed to SOS conditions for dynamics with polynomial or rational vector fields. Although computationally tractable, it still has some limitations in that the problem size grows exponentially with the number of variables and basis functions Tan [2006]. Learning-based and data-driven control using contraction theory Tsukamoto and Chung [2021], Tsukamoto et al. [2021, 2021], Tsukamoto and Chung [2021] has been developed to refine these ideas, using the high representational power of DNNs Hornik et al. [1989], Cybenko [1989], Funahashi [1989] and their scalable training realized by stochastic gradient descent Robbins and Monro [1951], Wensing and Slotine [2020].
The major advantage of using contraction theory for learning-based and data-driven control is that, by regarding its internal learning error as an external disturbance, we can ensure the distance between the target and learned trajectories to be bounded exponentially with time as in the CV-STEM results Tsukamoto and Chung [2021], Tsukamoto et al. [2021, 2021], Tsukamoto and Chung [2021], with its steady-state upper bound proportional to the learning error. Such robustness and incremental stability guarantees are useful for formally evaluating the performance of machine learning techniques such as reinforcement learning Sutton and Barto [1998], Bertsekas and Tsitsiklis [1996], Everett et al. [2018], Berkenkamp et al. [2017], imitation learning Carius et al. [2020], Rivière et al. [2020], Ho and Ermon [2016], Gupta et al. [2018], Kuefler et al. [2017], or neural networks Hornik et al. [1989], Cybenko [1989], Funahashi [1989]. This implies contraction theory could be utilized as a central tool in realizing safe and robust operations of learning-based and data-driven control, estimation, and motion planning schemes in real-world scenarios. We especially focus on the following areas of research.
@@
In order to achieve real-time computation of a contraction metric, mathematical models based on a Deep Neural Network (DNN) called a Neural Contraction Metric (NCM) Tsukamoto and Chung [2021] and Neural Stochastic Contraction Metric (NSCM) Tsukamoto et al. [2021] are derived to compute optimal CV-STEM contraction metrics for nonlinear systems perturbed by deterministic and stochastic disturbances, respectively. It can be proven that the NCM and NSCM still yield robustness and optimality associated with the CV-STEM framework despite having non-zero modeling errors Tsukamoto et al. [2021]. These metrics could also be synthesized and learned simultaneously with their feedback control laws directly by DNNs Sun et al. [2020, 2021], Tsukamoto et al. [2021] at the expense of the convex property in the CV-STEM formulation.
@@
In Singh et al. [2017], Zhao et al. [2021], contraction theory is leveraged to develop a tracking feedback controller with an optimized control invariant tube, solving the problem of robust motion planning under bounded external disturbances. This problem is also considered for systems with changing operating conditions Lopez et al. [2019] and parametric uncertainty Lakshmanan et al. [2020], Gahlawat et al. [2021], Sun et al. [2021] for its broader use in practice. As these methods still require online computation of a target trajectory, Learning-based Autonomous Guidance with RObustness and Stability (LAG-ROS) Tsukamoto and Chung [2021] is developed to model such robust control laws including the CV-STEM by a DNN, without explicitly requiring the target or desired trajectory as its input. While this considers motion planning algorithms only implicitly to avoid solving them in real-time, it is shown that contraction theory still allows us to assure a property of robustness against deterministic and stochastic disturbances following the same argument as in the NCM and NSCM work Tsukamoto and Chung [2021]. Note that LAG-ROS using contraction theory is not intended to derive new learning-based motion planning, but rather to augment any existing motion planner with a real-time method of guaranteeing formal incremental robustness and stability, and thus still applicable to other methods such as tube-based robust Model Predictive Control (MPC) Mayne et al. [2011, 2005], Singh et al. [2017], Bemporad and Morari [1999], Zhao et al. [2021], Hewing et al. [2020], its dynamic and adaptive counterparts Lopez et al. [2019], Lakshmanan et al. [2020], Gahlawat et al. [2021], Sun et al. [2021], and CCM-based learning certified control Sun et al. [2020].
@@
Adaptive control using contraction theory is studied in Sharma and Kar [2009] for parametric strict-feedback nonlinear systems, and recently generalized to deal with systems with unmatched parametric uncertainty by means of parameter-dependent CCM feedback control Lopez and Slotine [2021, 2021]. This method is further explored to develop an adaptive Neural Contraction Metric (aNCM) Tsukamoto et al. [2021], a parameter-dependent DNN model of the adaptive CV-STEM contraction metric. As the name suggests, the aNCM control makes adaptive control of Lopez and Slotine [2021, 2021] implementable in real-time for asymptotic stabilization, while maintaining the learning-based robustness and CV-STEM-type optimality of the NCM. Although it is designed to avoid the computation for evaluating integrals involving geodesics unlike Lopez and Slotine [2021, 2021], these differential state feedback schemes could still be considered, trading off added computational cost for generality. It is demonstrated in Tsukamoto et al. [2021] that the aNCM is applicable to many types of systems such as robotics systems [Slotine and Li, 1991, p. 392], spacecraft high-fidelity dynamics Battin [1999], Morgan et al. [2012], and systems modeled by basis function approximation and DNNs Nelles [2001], Sanner and Slotine [1992]. Discrete changes could be incorporated in this framework using Boffi et al. [2020], El Rifai et al. [2005].

### v8

@@ Abstract
1. Introduction
Lyapunov theory is one of the most widely-used approaches to stability analysis of a nonlinear system Isidori:1995:NCS:545735, marino1995nonlinear, Khalil:1173048, Vidyasagar, Ref_Slotine, Nijmeijer, which provides a condition for stability with respect to an equilibrium point, a target trajectory, or an invariant set. Contraction theory Ref:contraction1, Ref:contraction4, Ref:contraction5, Ref:contraction2, Ref:contraction3 rewrites suitable Lyapunov stability conditions using a quadratic Lyapunov function of the differential states, defined by a Riemannian contraction metric and its uniformly positive definite matrix, thereby characterizing a necessary and sufficient condition for incremental exponential convergence of the multiple nonlinear system trajectories to one single trajectory. It can be regarded as a generalization of Krasovskii’s theorem [Ref_Slotine, p. 83] applied to nonlinear incremental stability analysis Ref:contraction1, 989067, where the differential formulation permits a pure differential coordinate change with a non-constant metric for simplifying its stability proofs Ref:contraction1.
The differential nature of contraction theory implies we can exploit the Linear Time-Varying (LTV) systems-type techniques for nonlinear stability analysis and control/estimation synthesis mypaperTAC, ncm, nscm, Ref:Stochastic, mypaper (see Table 1). We emphasize that some of these methodological simplifications in contraction theory are accomplished by its extensive use of exponential stability along with the comparison lemma [Khalil:1173048, pp. 102-103, pp. 350-353], in lieu of Input-to-State Stability (ISS) or uniform asymptotic stability which often renders nonlinear stability analysis more involved Isidori:1995:NCS:545735, marino1995nonlinear, Khalil:1173048, Vidyasagar, Ref_Slotine, Nijmeijer. Several studies related to the notion of contraction, although not based on direct differential analysis, can be traced back to Hart, Demi, Lew.
@@ 1. Introduction
7. $\mathcal{L}_{2}$ stability condition Reduces to LMI conditions in terms of contraction metric defined by positive definite matrix $M$ Hamilton-Jacobi inequality (PDE) in terms of Lyapunov function $V$ [Khalil:1173048, p. 211]
8. Modular stability Differential analysis handles hierarchical, feedback, and parallel combinations Ref:contraction2 Passivity is not intuitive for hierarchical combinations
@@ 1.1. Paper Organization
In Sec. 2, we present fundamental results of contraction theory for nonlinear robustness and stability analysis. Sections 3 and 4 consider nonlinear optimal feedback control and estimation problems from the perspective of contraction theory, deriving and delineating a convex optimization-based method for constructing contraction metrics. Section 3 also presents some new results on relating contraction theory to the bounded real lemma lmi and Kalman–Yakubovich–Popov (KYP) lemma [kypbook, p. 218].
@@ 1.2. Related Work
According to contraction theory, all the solution trajectories of a given nonlinear system converge to one single trajectory incrementally and exponentially, regardless of the initial conditions, if the system has a contraction metric and its associated quadratic Lyapunov function of the differential state Ref:contraction1 (see 1133 of Table 1). This paper primarily considers this generalized notion of stability, called incremental exponential stability, which enables systematic learning-based and data-driven control synthesis with formal robustness and stability guarantees. The purpose of this section is not for proposing that other notions of stability, such as traditional Lyapunov-based stability or incremental asymptotic stability for semi-contraction systems, should be replaced by incremental exponential stability, but for clarifying its advantages to help determine which of these concepts is the best fit when analyzing nonlinear robustness and stability (see Sec. 2 for the illustrative examples).
@@
In keeping with the use of the comparison lemma [Khalil:1173048, pp. 102-103, pp. 350-353], incremental exponential stability naturally holds for non-autonomous nonlinear systems without any additional conditions or modifications unlike Lyapunov techniques (see, e.g., the examples and theorems in Rantzer). Such aspects of contraction theory, including the extensive use of exponential stability, result in intuitive proofs on ISS and finite-gain $\mathcal{L}_{p}$ stability both for autonomous and non-autonomous nonlinear systems, without resorting to uniform asymptotic stability which makes stability analysis much more involved than necessary Isidori:1995:NCS:545735, marino1995nonlinear, Khalil:1173048, Vidyasagar, Ref_Slotine, Nijmeijer. In particular, perturbed systems with a time-varying target trajectory are non-autonomous, and thus contraction theory allows us to easily obtain an explicit exponential bound on its tracking error Ref:contraction1, Ref:phasesync, mypaperTAC, Ref:Stochastic, mypaper, leveraging incremental stability of the perturbed system trajectories with respect to the target trajectory (see $4$ of Table 1). Having such analytical bounds on the tracking error is almost essential for the safe and robust implementation of automatic control schemes in real-world scenarios.
@@
Contraction theory also simplifies input-output stability analysis, such as $\mathcal{L}_{p}$ gain analysis of nonlinear systems including the $\mathcal{H}_{\infty}$ nonlinear optimal control problem 256331, 159566, 29425, 85062, 151101, 599969, doi:10.1137/S0363012996301336, 1259458. For example, in Lyapunov theory, the problem of finding a suitable Lyapunov function with the smallest $\mathcal{L}_{2}$ gain boils down to solving a Partial Differential Equation (PDE) called the Hamilton-Jacobi inequality [Khalil:1173048, p. 211]256331 in terms of its associated Lyapunov function. In essence, since contraction theory utilizes a quadratic Lyapunov function of the differential state for stability analysis, the problem could be solved with a Linear Matrix Inequality (LMI) constraint lmi analogous to the KYP lemma [kypbook, p. 218] in LTV systems theory mypaperTAC, Ref:Stochastic, mypaper, rccm, as shall be shown in Sec. 34 (see 5577 of Table 1). There exist stochastic analogues of these stability results for nonlinear systems with stochastic perturbations Pham2009, mypaperTAC, Ref:Stochastic, mypaper, han2021incremental, as shall be outlined also in this paper.
@@
Another notable feature of contraction theory is modularity, which preserves contraction through parallel, feedback, and hierarchical combinations Ref:contraction2, Ref:contraction5, specific time-delayed feedback communications 1618853, synchronized coupled oscillations Ref:contraction3, Ref:contraction_robot, and synchronized networks Ref:contraction2, Ref:phasesync, Ref:contraction_sync, Ref:contraction_robot, 6327337, Ref:phasesync, Ref:ChungTRO, expanding the results obtainable with the passivity formalism of Lyapunov theory passivity[Khalil:1173048, p. 227][Ref_Slotine, p. 132] (see $8$ of Table 1). Due to all these useful properties, extensions of contraction theory have been considered in many different settings. These include, but are not limited to, stochastic contraction (Gaussian white noise Pham2009, mypaperTAC, Ref:Stochastic, mypaper, Poisson shot noise and Lévy noise han2021incremental), contraction for discrete and hybrid nonlinear systems Ref:contraction1, Ref:contraction_robot, Ref:contraction4, 4795665, mypaperTAC, mypaper, wei2021discretetime, partial contraction Ref:contraction3, transverse contraction MANCHESTER201432, incremental stability analysis of nonlinear estimation (the Extended Kalman Filter (EKF) 6849943, nonlinear observers zhao2005discrete, Ref:Stochastic, Simultaneous Localization And Mapping (SLAM) doi:10.1177/0278364917710541), generalized gradient descent based on geodesical convexity beyondconvexity, contraction on Finsler and Riemannian manifolds 6632882, scrm, SIMPSONPORCO201474, finsler_ccm, contraction on Banach and Hilbert spaces for PDEs 6580717, 6161268, pdecontraction, non-Euclidean contraction sontag_contraction, davydov2021noneuclidean, contracting learning with piecewise-linear basis functions lohmiller2018notes, incremental quadratic stability analysis quad_stability, contraction after small transients MARGALIOT2016178, immersion and invariance stabilizing controller design 7581040, 1193738, and Lipschitz-bounded neural networks for robustness and stability guarantees neurallander, revay2020lipschitz, revay2021recurrent, cdc_systemid.
@@
The benefits of contraction theory reviewed so far naturally lead to a discussion on how to design a contraction metric and corresponding Lyapunov function. There are some cases in which we can analytically find them using special structures of systems in question Isidori:1995:NCS:545735, Ref_Slotine, Khalil:1173048. Among these are Lagrangian systems [Ref_Slotine, p. 392], where one easy choice of positive definite matrices that define a contraction metric is the inertia matrix, or feedback linearizable systems 532343, Primbs99nonlinearoptimal, 1425952, AYLWARD20082163, 5443730, where we could solve the Riccati equation for a contraction metric as in LTV systems. This is also the case in the context of state estimation (e.g., the nonlinear SLAM problem can be reformulated as an LTV estimation problem using virtual synthetic measurements doi:10.1177/0278364917710541, Ref:Stochastic). Once we find a contraction metric and Lyapunov function of a nominal nonlinear system for the sake of stability, they could be used as a Control Lyapunov Function (CLF) to attain stabilizing feedback control doi:10.1137/0321028, SONTAG1989117, Khalil:1173048 or could be augmented with an integral control law called adaptive backstepping to recursively design a Lyapunov function for strict- and output-feedback systems backstepping, DENG1997143, Deng1999, 5160111. However, deriving an analytical form of contraction metrics for general nonlinear systems is challenging, and thus several search algorithms have been developed for finding them at least numerically using the LMI nature of the contraction condition.
@@
The simplest of these techniques is the method of State-Dependent Riccati Equation (SDRE) sddre, Banks2007, survey_SDRE, which uses the State-Dependent Coefficient (SDC) parameterization (also known as extended linearization) of nonlinear systems for feedback control and state estimation synthesis. Motivating optimization-based approaches to design a contraction metric, it is proposed in mypaperTAC, Ref:Stochastic, mypaper that the Hamilton-Jacobi inequality for the finite-gain $\mathcal{L}_{2}$ stability condition can be expressed as an LMI when contraction theory is equipped with the extended linearity of the SDC formulation. Specifically, in mypaperTAC, ncm, nscm, mypaper, a convex optimization-based framework for robust feedback control and state estimation, named ConVex optimization-based Steady-state Tracking Error Minimization (CV-STEM), is derived to find a contraction metric that minimizes an upper bound of the steady-state distance between perturbed and unperturbed system trajectories. In this context, we could utilize Control Contraction Metrics (CCMs) ccm, 7989693, 47710, WANG201944, vccm, rccm for extending contraction theory to the systematic design of differential feedback control $\delta u=k(x,\delta x,u,t)$ via convex optimization, achieving greater generality at the expense of computational efficiency in obtaining $u$. Applications of the CCM to estimation, adaptive control, and motion planning are discussed in estimation_ccm9109296, lopez2021universal, regretboffi, and 7989693, 8814758, 9303957, L1contraction2, sun2021uncertaintyaware, respectively, using geodesic distances between trajectories scrm. It is also worth noting that the objective function of CV-STEM has the condition number of a positive definite matrix that defines a contraction metric as one of its arguments, rendering it applicable and effective even to machine learning-based automatic control frameworks as shall be seen in Sec. 59.
@@
One drawback of these numerical schemes is that they require solving optimization problems or nonlinear systems of equations at each time instant online, which is not necessarily realistic in practice. In Lyapunov theory, approximating functions in a given hypothesis space has therefore been a standard technique 664157, JOHANSEN20001617, LyapunovRBF, 1184414, spencer18lyapunovnn, NIPS2019_8587, 8263816, 9115021, 9146356, 9147615, gaby2021lyapunovnet, where examples of its function classes include piecewise quadratic functions 664157, linearly parameterized non-quadratic functions JOHANSEN20001617, a linear combination of radial basis functions LyapunovRBF, Sum-Of-Squares (SOS) functions 1184414, and neural networks spencer18lyapunovnn, NIPS2019_8587, gaby2021lyapunovnet. In AYLWARD20082163, ccm, the SOS approximation is investigated for the case of contraction theory, showing that the contraction condition can be relaxed to SOS constraints for dynamics with polynomial or rational vector fields. Although computationally tractable, it still has some limitations in that the problem size grows exponentially with the number of variables and basis functions sos_dissertation. Learning-based and data-driven control using contraction theory ncm, nscm, ancm, lagros has been developed to refine these ideas, using the high representational power of DNNs neural1, neural2, neural3 and their scalable training realized by stochastic gradient descent sgd, beyondconvexity.
The major advantage of using contraction theory for learning-based and data-driven control is that, by regarding its internal learning error as an external disturbance, we can ensure the distance between the target and learned trajectories to be bounded exponentially with time as in the CV-STEM results ncm, nscm, ancm, lagros, with its steady-state upper bound proportional to the learning error. Such robustness and incremental stability guarantees are useful for formally evaluating the performance of machine learning techniques such as reinforcement learning sutton, ndp, 8593871, NIPS2017_766ebcd5, imitation learning 9001182, glas, NIPS2016_cc7e2b87, 8578338, 7995721, or neural networks neural1, neural2, neural3. This implies contraction theory could be utilized as a central tool in realizing safe and robust operations of learning-based and data-driven control, estimation, and motion planning schemes in real-world scenarios. We especially focus on the following areas of research.
@@
In order to achieve real-time computation of a contraction metric, mathematical models based on a Deep Neural Network (DNN) called a Neural Contraction Metric (NCM) ncm and Neural Stochastic Contraction Metric (NSCM) nscm are derived to compute optimal CV-STEM contraction metrics for nonlinear systems perturbed by deterministic and stochastic disturbances, respectively. It can be proven that the NCM and NSCM still yield robustness and optimality associated with the CV-STEM framework despite having non-zero modeling errors cdc_ncm. These metrics could also be synthesized and learned simultaneously with their feedback control laws directly by DNNs chuchu, sun2021uncertaintyaware, cdc_ncm at the expense of the convex property in the CV-STEM formulation.
@@
In 7989693, zhao2021tubecertified, contraction theory is leveraged to develop a tracking feedback controller with an optimized control invariant tube, solving the problem of robust motion planning under bounded external disturbances. This problem is also considered for systems with changing operating conditions 8814758 and parametric uncertainty 9303957, L1contraction2, sun2021uncertaintyaware for its broader use in practice. As these methods still require online computation of a target trajectory, Learning-based Autonomous Guidance with RObustness and Stability (LAG-ROS) lagros is developed to model such robust control laws including the CV-STEM by a DNN, without explicitly requiring the target or desired trajectory as its input. While this considers motion planning algorithms only implicitly to avoid solving them in real-time, it is shown that contraction theory still allows us to assure a property of robustness against deterministic and stochastic disturbances following the same argument as in the NCM and NSCM work lagros. Note that LAG-ROS using contraction theory is not intended to derive new learning-based motion planning, but rather to augment any existing motion planner with a real-time method of guaranteeing formal incremental robustness and stability, and thus still applicable to other methods such as tube-based robust Model Predictive Control (MPC) tube_nmpc, tube_mpc, 7989693, 10.1007/BFb0109870, zhao2021tubecertified, learningmpc, its dynamic and adaptive counterparts 8814758, 9303957, L1contraction2, sun2021uncertaintyaware, and CCM-based learning certified control chuchu.
@@
Adaptive control using contraction theory is studied in 5160111 for parametric strict-feedback nonlinear systems, and recently generalized to deal with systems with unmatched parametric uncertainty by means of parameter-dependent CCM feedback control 9109296, lopez2021universal. This method is further explored to develop an adaptive Neural Contraction Metric (aNCM) ancm, a parameter-dependent DNN model of the adaptive CV-STEM contraction metric. As the name suggests, the aNCM control makes adaptive control of 9109296, lopez2021universal implementable in real-time for asymptotic stabilization, while maintaining the learning-based robustness and CV-STEM-type optimality of the NCM. Although it is designed to avoid the computation for evaluating integrals involving geodesics unlike 9109296, lopez2021universal, these differential state feedback schemes could still be considered, trading off added computational cost for generality. It is demonstrated in ancm that the aNCM is applicable to many types of systems such as robotics systems [Ref_Slotine, p. 392], spacecraft high-fidelity dynamics battin, doi:10.2514/1.55705, and systems modeled by basis function approximation and DNNs Nelles2001, SannerSlotine1992. Discrete changes could be incorporated in this framework using regretboffi, 1469901.